OpenStudy (anonymous):

(x + y)(x^2 - xy + y^2) helppp:(

OpenStudy (gw2011):

Multiply each item as follows (multiply each item in the left-hand side with each item in the right-hand): (x)(x^2)=x^3 (x)(-xy)=-x^2y (x)(y^2)=xy^2 (y)(x^2)=yx^2=x^2y (y)(-xy)=-xy^2 (y)(y^2)=y^3 Now combining this you get: x^3-x^2y+xy^2+x^2y-xy^2+y^3=x^3+y^3 The answer is: x^3+y^3